OpenStudy (kkutie7):

Getting a little confused on this problem. A function has the following Taylor series: I'll post the series and my work in few.

OpenStudy (kkutie7):

\[f(x)=\sum_{k=0}^{\infty}(-1)^{k+1}\frac{ k! }{ (2k)! }(x-4)^{k}\]

OpenStudy (kkutie7):

find (a) f(4) (b) f'(4) (c) f"(4)

OpenStudy (kkutie7):

I wrote out the first 3 non zero terms: \[-1+\frac{ x-4 }{ 2 }-\frac{ (x-4)^{2} }{ 24 }\]

OpenStudy (kkutie7):

I they took a shot at the first one like this (Keep in mind I have no clue if I'm even remotely doing this correct. Probably not.) \[\frac{ f^{0}(0)x^{0} }{ 0! }=-1\rightarrow f^{0}(0)=\frac{ -1*0! }{ x^{0} }\rightarrow f^{0}(0)=-1\] For the other two I did the same thing and got zero. Zero is of course wrong.

OpenStudy (anonymous):

Just remember that a Taylor series is given by \[ f(x) = \sum_n a_n(x-x_0)^n = \sum_n \frac{f^{(n)}(x_0)}{n!} (x-x_0)^n\] So if you know the terms of the series, you should be able to simply read off the derivatives: \[f(x) = \sum_k(-1)^{k+1} \frac{k!}{(2k)!}(x-4)^k = \sum_k \frac{f^{(k)}(4)}{k!}(x-4)^k\] which implies that \[f^{(k)}(4) = (-1)^{k+1} \frac{k!k!}{(2k)!} \]

OpenStudy (anonymous):

so \[k=0 \rightarrow f(4) = -1\] \[k=1 \rightarrow f'(4) = \frac{1}{2} \] \[k=2 \rightarrow f''(4) = -\frac{2!2!}{4!} = -\frac{4}{4!} = -\frac{1}{6} \] As a side note, the denominator in the third term in the expansion should be 12, not 24.